# Optimal Sizing and Material Choice for Additively Manufactured Compact Plate Heat Exchangers

### v1

@@ 3 Material selection for maximal power density
A wide variety of materials prove suitable for the manufacture of heat exchangers, and the choice is typically driven by the requirements and operating conditions of the device. Stainless steel, for example, finds widespread applications in heat exchanger manufacturing due to its resistance to corrosion and high temperatures [33]. Copper is often preferred in heat transfer applications due to its superior thermal conductivity [4]. Aluminum stands out as a common choice due to its lightweight, excellent thermal conductivity, and affordability compared to other metals [34]. Ceramic materials are increasingly considered advantageous for heat exchangers due to their temperature and corrosion resistance, outperforming traditional metallic variants [35]. Beyond metals and ceramics, polymers also offer advantages in heat exchanger manufacturing, such as cost-effectiveness, lightweight construction, and resistance to corrosion and fouling [9].
[…]
First, we investigate the effect of different materials on the maximum power density at a given effectiveness, without accounting for manufacturing and fouling constraints. To achieve this, we formulate an optimization problem that investigates the maximal power density that can be achieved for each material as a function of the Bejan number Be, and the relative plate spacing $\widetilde{D}$. This optimization problem is defined as follows:
$\max_{\widetilde{D}\in\mathbb{R}_{0}^{+},\,\textup{Be}\in\mathbb{R}_{0}^{+}}$ $\qquad Q^{\diamond}\left(\widetilde{D},\textup{Be}\right),$ (13)
s.t. $\qquad\varepsilon\left(\widetilde{D},\textup{Be}\right)=\varepsilon_{d},$
where $\varepsilon_{d}$ is the design effectiveness.
Rogiers and Baelmans [27] demonstrated that the optimal Bejan number $\textup{Be}_{opt}$, which maximizes power density, also maximizes effectiveness when the relative plate thickness $t/D$ is kept constant in the optimization problem. Their method simplifies the original two-dimensional optimization problem into two sequential steps. In the first step, for a given relative plate spacing $\widetilde{D}$, the optimal Bejan number $\textup{Be}^{*}$ is determined through the following one-dimensional unconstrained optimization problem:
$^{*}=\operatorname*{arg\,max}_{\textup{Be}\in\mathbb{R}_{0}^{+}}\varepsilon\left(\widetilde{D},\textup{Be}\right).$ Be ) . (14)
This equation yields the Bejan number that maximizes effectiveness for the specified $\widetilde{D}$. However, this intermediate result does not fully solve the optimization problem, as the optimal relative plate spacing, $\widetilde{D}_{opt}$, remains to be determined. In the second step, $\widetilde{D}_{opt}$ is identified by ensuring that the corresponding maximum effectiveness meets the design requirement:
$\varepsilon_{max}\left(\widetilde{D}\right)=\varepsilon_{d}.$ (15)
The global optimal Bejan number is then obtained as $\textup{Be}_{opt}=\textup{Be}^{*}(\widetilde{D}_{opt})$. The resulting optimal variables, $\textup{Be}_{opt}$ and $\widetilde{D}_{opt}$, represent the Bejan number and relative plate spacing of the heat exchanger that achieves the maximum power density while satisfying the design effectiveness.
To optimize a cHEX design using various materials, this simplified method is applied with a design effectiveness $\varepsilon_{d}$ of 0.79, corresponding to the effectiveness of the initial heat exchanger, detailed in Table 1, that is used as the baseline. To maintain consistency with the baseline geometry, the constant $t/D$ is fixed at 0.16, and the target pressure drop for the optimization is set equal to the pressure drop of the initial design in Table 1. Since the thermal model assumes constant fluid properties, the thermophysical properties of the working fluid are evaluated at the average inlet temperature of the cold and hot fluid channels in the baseline case. These properties are also listed in Table 1.
Table 1: Working fluid and initial heat exchanger properties
@@
[…]
Effectiveness, $\varepsilon$ 0.79
Material Austenitic steel
The optimization results are converted into dimensional parameters, and the optimal geometrical parameters, together with the corresponding power densities, $\mathcal{D}$, for each heat exchanger made of different materials, are presented in Table 2. Furthermore, the table provides a comparative analysis of these optimized heat exchangers against the baseline case, referred to as “austenitic steel (initial).” This comparison is expressed as the ratio of the power density of the optimal configuration to that of the initial heat exchanger, defined as the Improvement Factor, IF.
Table 2: Dimensional properties of initial and optimized heat exchangers without manufacturing constraints, with associated improvement factors
[…]
In the optimal configurations, the geometrical parameters (length, spacing, and thickness) increase with the material’s thermal conductivity, leading to a reduction in power density, $\mathcal{D}$. For instance, in the plastic heat exchanger, which has the lowest plate conductivity, $\mathcal{D}$ increases by approximately 1785 times compared to the initial heat exchanger. In contrast, the copper heat exchanger, with the highest plate conductivity, experiences a slight decrease in $\mathcal{D}$ by a factor of 0.98, making copper the only material for which $\mathcal{D}$ is lower than in the initial geometry.
A comparison of the copper heat exchanger’s geometrical parameters with the baseline reveals that its length, spacing, and thickness are all smaller. However, despite this reduction in size, the lower power density is attributed to a decrease in the mass flow rate per unit width for a single channel $\dot{m}^{\prime}_{\text{single}}$, caused by the smaller channels. To compensate for this reduced $\dot{m}^{\prime}_{\text{single}}$, the number of layers $n$, the width $W$, or both must be increased, resulting in a larger total volume to achieve the same thermal power as the initial heat exchanger. To ensure that the optimal heat exchanger delivers the same thermal power with the same effectiveness as the initial configuration, $n$ and $W$ must be adjusted such that $(nW\dot{m}^{\prime}_{\text{single}})_{\text{initial}}=(nW\dot{m}^{\prime}_{\text{single}})_{\text{optimal}}.$
Without any constraints on plate thickness and spacing, the optimal configurations feature plate thicknesses ranging from 0.0025 mm to 0.11 mm. However, such thin structures raise concerns about structural integrity and manufacturability. The AM process cannot reliably ensure the producibility or reproducibility of these thin features, and the accuracy error increases as the component thickness decreases [4]. This challenge is especially pronounced for copper and its alloys due to their high surface reflectivity and low laser absorption. Additionally, the plate spacings in the optimal heat exchangers range from 0.016 mm to 0.7 mm, increasing the risk of fouling. Therefore, the optimization problem must be revised to address these limitations. The next section presents the formulation of this revised optimization problem and the results of the new approach.
@@ 4 Design optimization considering additive manufacturing and fouling constraints
The goal of this section is to design a cHEX with minimal size while maintaining a given pressure drop and effectiveness, considering constraints related to manufacturing, fouling, and structural limitations on plate thickness and spacing. In cHEX design, thinner plates are generally preferred because they reduce axial heat conduction and minimize lateral conduction resistance, enhancing thermal performance. However, the minimum feasible plate thickness is influenced by the selected manufacturing process and material type. Therefore, it should be explicitly included in the optimization process as a design constraint.
In the previous section, where the relative plate thickness $t/D$ was kept constant, the optimization problem could be simplified to a one-dimensional, unconstrained optimization, since the Bejan number maximized power density also maximized the effectiveness. This simplification no longer applies in the current section, as the points of maximum effectiveness and maximum power density differ when using a specified design plate thickness $t_{d}$ instead of a fixed relative plate thickness $t/D$. Figure 2 illustrates this distinction by showing the variation in effectiveness and non-dimensional power density as a function of the Bejan number, for a given $\widetilde{D}=0.0106$ and $t_{d}=0.5mm$. Hence, the original problem definition in Eq. 13 is altered to obtain the following optimization problem:
$\max_{\widetilde{D}\in\mathbb{R}_{0}^{+},\,\textup{Be}\in\mathbb{R}_{0}^{+}}$ $\qquad Q^{\diamond}\left(\widetilde{D},\textup{Be}\right),$ (16)
$s.t.$ $\qquad\varepsilon\left(\widetilde{D},\textup{Be}\right)=\varepsilon_{d},$
$\qquad t=t_{d},$
$\qquad D\geq D_{min}.$
In addition to the plate thickness, a restriction on the plate spacing is imposed to reduce the risk of fouling. To mitigate this risk, the additional constraint ensures that the plate spacing $D$ is not less than a specified minimum value $D_{min}$. This formulation results in a constrained nonlinear optimization problem. The problem is solved numerically using Python’s SciPy optimization library, specifically employing the minimize function, which supports various constrained optimization algorithms [36]. In this study, Sequential Least Squares Programming (SLSQP) was used to obtain the optimal values of $\widetilde{D}$ and Be while satisfying the given constraints.
Figure 2: Effectiveness and non-dimensional power density profiles with Bejan Number for $\widetilde{D}=0.0106$ and $t=0.5mm$
@@
The optimizations in this section use the same fluid properties, pressure drop, and effectiveness as the baseline heat exchanger, outlined in the previous section and detailed in Table 1. The results of these optimizations are illustrated in Figure 3, which compares three distinct optimization strategies aimed at maximizing power density.
@@
The first strategy, labeled “Thickness Constraint”, maintains a fixed $t_{d}$=0.5 mm for all materials, with no constraint on plate spacing. The second strategy, labeled “Thickness and Fouling Constraints”, introduces a fouling constraint that imposes $D_{min}$=0.8 mm, while still maintaining the fixed plate thickness of 0.5 mm for all materials. The third strategy, labeled “Material-Specific Thickness and Fouling Constraints”, incorporates both plate thickness and fouling constraints simultaneously. Unlike the previous strategies, this approach applies a material-specific thickness constraint to account for variations in manufacturing resolution in AM.
@@
To address AM resolutions for ceramic and metal heat exchangers, the plate thickness is determined based on the melt pool size, a characteristic of the manufacturing process. It is assumed that a sealed plate requires a thickness of twice the melt pool size. Consequently, the following plate thickness values are used: 0.5 mm for copper [37], 0.3 mm for aluminum [38], and 0.25 mm for aluminum nitride (AlN), aluminum oxide (Al2O3[39], and austenitic steel [40]. These values are characteristic of the Laser Powder Bed Fusion (LPBF) process, which is the most common technique for AM of metals and ceramics. For plastic heat exchangers, the plate thickness is not derived from melt pool size, as LPBF is not typically used for plastic printing. Instead, Material Extrusion (MEX), Vat Photopolymerization (VPP), and Sheet Lamination (SHL) processes are widely used [41, 42]. While these techniques can achieve thin walls with resolutions below 0.1 mm, a plate thickness of 0.1 mm is assumed in this study. This value is selected to balance manufacturing feasibility with structural integrity, as heat exchangers with wall thicknesses below 0.1 mm are prone to plastic deformation [43].
In addition to the three optimization strategies, Figure 3 presents the non-dimensional power density values for heat exchangers with baseline length and plate spacing, labeled as “Additive Manufacturable Initial” in the figure. These designs serve as a reference for comparison and are adapted from the baseline heat exchanger by applying material-specific plate thicknesses for each of the six considered materials. Unlike in the optimal designs, the pressure drop in these cases is lower than 1150 Pa to maintain a constant effectiveness of 0.79. As a result, $\psi$ in the non-dimensional power density equation (Eq. 5) is no longer equal to 1; instead, it takes values of 0.96, 0.99, 0.98, and 0.75 for plastic, austenitic steel, AlN, and Al2O3, respectively. However, for copper and aluminum, even with adjustments to the pressure drop, the design with the baseline length and plate spacing is insufficient to achieve the target effectiveness. Therefore, no Additive Manufacurable Initial designs are included in the figure for these materials.
@@
Figure 3: Maximum non-dimensional power densities obtained from optimizations with $\varepsilon_{d}~{}=~{}0.79$
@@
To provide additional context for the designs presented in Figure 3, Table 3 lists the key geometrical dimensions of the heat exchangers for each optimization strategy and the reference designs. The table includes values such as plate thickness, plate spacing, and overall heat exchanger length, reflecting the influence of material properties and manufacturing constraints on the final optimized geometries.
Table 3: Geometrical dimensions of heat exchanger designs for different optimization strategies and reference cases with $\varepsilon_{d}=0.79$
Case Material t [mm] D [mm] L [mm]
Thickness Constraint Plastic 0.5 0.14 5.6
Austenitic steel 0.5 0.48 46.0
Ceramic (Al2O3) 0.5 0.54 56.8
Ceramic (AlN) 0.5 0.95 188.2
Aluminum 0.5 1.03 223.6
Copper 0.5 1.19 310.4
Thickness and Fouling Constraints Plastic 0.5 0.8* 116.4
Austenitic steel 0.5 0.8* 105.3
Ceramic (Al2O3) 0.5 0.8* 107.2
Ceramic (AlN) 0.5 0.95 188.2
Aluminum 0.5 1.03 223.6
Copper 0.5 1.19 310.4
Material-specific Thickness and Fouling Constraints Plastic 0.1 0.8* 103.8
Austenitic steel 0.25 0.8* 102.8
Ceramic (Al2O3) 0.25 0.8* 103.7
Ceramic (AlN) 0.25 0.8* 126.6
Aluminum 0.3 0.88 161.6
Copper 0.5 1.19 310.4
Additive Manufacturable Initial Plastic 0.1 1.00 158.0
Austenitic steel 0.25 1.00 158.0
Ceramic (Al2O3) 0.25 1.00 158.0
Ceramic (AlN) 0.25 1.00 158.0
@@
As shown in Figure 3, under the “Thickness Constraint” strategy, where the plate thickness is fixed at 0.5 mm and the fouling constraint is not considered, plastic, the material with the lowest thermal conductivity, achieves the highest power density. This result corresponds to a plate spacing of 0.14 mm and a length of 5.6 mm, dimensions that are significantly smaller than those of all other designs (see Table 3). When the fouling constraint is included, austenitic steel emerges as the material with the highest power density, while the plastic heat exchanger drops to third place, following Al2O3. However, in the “Material-Specific Thickness and Fouling Constraints” strategy, where material-specific resolutions are explicitly accounted for, plastic achieves the highest power density due to its ability to utilize thinner plates.
In contrast, despite having the highest thermal conductivity among the materials considered, copper exhibits the lowest power density across all optimization strategies in Figure 3. This indicates that the dimensions of copper heat exchangers need to be significantly higher than the others to achieve the design effectiveness of 0.79, as confirmed by Table 3.
Across all three optimization scenarios, heat exchangers made from relatively higher-conductivity materials such as copper and aluminum consistently result in increased length and spacing compared to those made from materials with lower thermal conductivity. Consequently, the fouling constraint for copper and aluminum remains inactive. As marked in Table 3, for AlN, the fouling constraint is inactive under the “Thickness and Fouling Constraints” strategy, where the plate thickness is 0.5 mm, but becomes active under the “Material-Specific Thickness and Fouling Constraints strategy, where the plate thickness is reduced to 0.25 mm. These results underscore that using thicker plates and high-conductivity materials in cHEX designs leads to larger heat exchangers to offset axial wall conduction effects and maintain the target design effectiveness.

### v2

@@ 3 Material selection for maximal power density
A wide variety of materials prove suitable for the manufacturing of heat exchangers, and the choice is typically driven by the requirements and operating conditions of the device. Stainless steel, for example, finds widespread applications in heat exchanger manufacturing due to its resistance to corrosion and high temperatures ZadiMaad2018 . Copper is often preferred in heat transfer applications due to its superior thermal conductivity Careri2023 . Aluminum stands out as a common choice due to its lightweight, excellent thermal conductivity, and affordability compared to other metals Cantor2015 . Ceramic materials are increasingly considered advantageous for heat exchangers due to their temperature and corrosion resistance, outperforming traditional metallic variants Sommers2010 . Beyond metals and ceramics, polymers also offer advantages in heat exchanger manufacturing, such as cost-effectiveness, lightweight construction, and resistance to corrosion and fouling Hein2021 . Considering this diversity, the current study investigates the effect of six different materials on the maximum power density. These materials are copper (with a thermal conductivity, $k_{\text{Cu}}=398~\text{W/m}\cdot\text{K}$), aluminum ($k_{\text{Al}}=237~\text{W/m}\cdot\text{K}$), aluminum nitride (AlN, $k_{\text{AlN}}=180~\text{W/m}\cdot\text{K}$), aluminum oxide (Al2O3, $k_{\text{Al\textsubscript{2}O\textsubscript{3}}}=27~\text{W/m}\cdot\text{K}$), austenitic steel ($k_{\text{A.st}}=20~\text{W/m}\cdot\text{K}$), and a typical plastic ($k_{\text{Pl}}=0.2~\text{W/m}\cdot\text{K}$).
First, we examine the effect of different materials on the maximum power density at a given effectiveness, without accounting for manufacturing and fouling constraints. To achieve this, we formulate an optimization problem that maximizes power density for each material, characterized by its thermal conductivity $k_{w}$, by varying plate spacing and channel length. To avoid unrealistic solutions where the optimizer reduces the wall thickness to zero, we impose a geometric constraint that links wall thickness in proportion to plate spacing: $t=\gamma D$, where $\gamma$ is a fixed constant. This prevents physically degenerate designs by ensuring the wall always retains finite thickness. The optimization problem is then written in dimensionless form as
$\max_{D^{\ast}>0,\ L^{\ast}>0,\ t^{\ast}>0}$ $\qquad Q^{\diamond}\left(D^{\ast},L^{\ast},t^{\ast}\right),$ (15)
s.t. $\qquad\varepsilon\left(D^{\ast},L^{\ast},t^{\ast};k_{w}/k\right)=\varepsilon_{d},$
$\qquad t^{\ast}=\gamma\cdot D^{\ast},$
where $\varepsilon_{d}$ is the design effectiveness. This formulation follows the same optimization structure introduced by Rogiers and Baelmans Rogiers2010 , who also maximized power density under a fixed effectiveness constraint. The nondimensional parameters used here differ, primarily due to the explicit treatment of wall thickness, which enables the integration of additional constraints in future design stages.
The nonlinear optimization problem is solved numerically for each of the six materials using Python’s SciPy optimization library, specifically the minimize function, which supports various constrained-solvers Virtanen2020 . In this study, the Sequential Least Squares Programming (SLSQP) algorithm is employed to compute the optimal values of $D^{\ast}$ and $L^{\ast}$, while satisfying the imposed constraints. This numerical procedure is used consistently for all optimization problems discussed in the paper. To ensure comparability with the baseline heat exchanger (Table 1), the design effectiveness is fixed at $\varepsilon_{d}=0.79$, and the proportionality constant is set to $\gamma=0.16$, corresponding to the original $t^{\ast}/D^{\ast}$ ratio. The thermophysical properties of the fluid, listed in Table 1, are evaluated at the mean of the inlet temperatures of the hot and cold fluid and are assumed constant throughout the analysis. The pressure drop is matched to the baseline to ensure a fair comparison across all cases.
Table 1: Working fluid and initial heat exchanger properties. The plate thickness and pressure drop listed here correspond to the reference values $t_{\text{ref}}$ and $\Delta P_{\text{ref}}$ used throughout the text.
@@
Fluid density, $\rho$ $1.060\text{\,}\mathrm{kg}\text{\,}{\mathrm{m}}^{-3}$
Fluid isobaric specific heat, $C_{p}$ $1.008\text{\,}\mathrm{kJ}\text{\,}{\mathrm{kg}}^{-1}\text{\,}{\mathrm{K}}^{-1}$
Fluid dynamic viscosity, $\mu$ $19.99\text{\times}{10}^{-6}\text{\,}\mathrm{Pa}\text{\,}\mathrm{s}$
Fluid thermal conductivity, $k$ $28.8\text{\,}\mathrm{mW}\text{\,}{\mathrm{m}}^{-1}\text{\,}{\mathrm{K}}^{-1}$
@@
Pressure drop, $\Delta P$ $170\text{\,}\mathrm{Pa}$
Effectiveness, $\varepsilon$ 0.79
Length, $L$ $158.0\text{\,}\mathrm{mm}$
Width, $W$ $95.0\text{\,}\mathrm{mm}$
Plate spacing, $D$ $1.0\text{\,}\mathrm{mm}$
Plate thickness, $t$ $0.16\text{\,}\mathrm{mm}$
Material Austenitic steel
The optimization results are summarized in Table 2, which reports the nondimensional geometrical parameters $L^{\ast}$, $D^{\ast}$, and $t^{\ast}$ that maximize power density for each material. The corresponding nondimensional power densities $Q^{\diamond}$ are also listed, along with the Improvement Factor (IF), defined as the ratio of the optimized power density to that of the baseline design, referred to as “austenitic steel (initial)”.
Table 2: Nondimensional properties of initial and optimized heat exchangers without manufacturing constraints, with associated improvement factors. All geometrical values are scaled by the reference wall thickness $t_{\text{ref}}=$0.16\text{\,}\mathrm{mm}$\$.
Material $k_{w}/k$ $L^{\ast}$ $D^{\ast}$ $t^{\ast}$ $Q^{\diamond}\times 10^{6}$ IF
Austenitic steel (Initial) $694.44$ 987.50 $6.250$ 1.000 $1.138$ 1.00
Plastic $6.94$ 0.70 $0.137$ 0.022 1 095.6151\,095.615 962.44
Austenitic steel $694.44$ 68.44 $1.375$ 0.220 $11.470$ 10.08
Ceramic (Al2O3) $937.50$ 91.51 $1.582$ 0.253 $8.495$ 7.46
Ceramic (AlN) 6 250.006\,250.00 615.53 $4.123$ 0.660 $1.275$ 1.12
Aluminum 8 229.178\,229.17 808.22 $4.718$ 0.755 $0.968$ 0.85
Copper 13 819.4413\,819.44 1361.25 $6.133$ 0.981 $0.577$ 0.51
In the optimal configurations, the nondimensional geometrical parameters $L^{\ast}$, $D^{\ast}$, and $t^{\ast}$ increase with the material’s thermal conductivity, leading to a reduction in nondimensional power density $Q^{\diamond}$. For instance, the plastic heat exchanger, which has the lowest plate conductivity, achieves the highest performance, with an IF of approximately 962 relative to the initial design. In contrast, materials with high conductivity, such as aluminum and copper, result in lower power densities than the baseline. Copper, in particular, exhibits the lowest performance, with an IF of 0.51, indicating a 49% reduction compared to the initial design.
To interpret the optimized nondimensional designs in physical terms, each configuration can be translated into dimensional quantities using the wall thickness of the initial heat exchanger configuration as the reference wall thickness $t_{\text{ref}}$. The nondimensional parameters $L^{\ast}$, $D^{\ast}$, and $t^{\ast}$ are rescaled to obtain the corresponding dimensional values $L$, $D$, and $t$. To maintain consistency with the baseline thermal power, the number of layers $n$ and the width $W$ are selected such that the product with the mass flow rate per unit width for a single channel, $\dot{m}^{\prime}_{\text{single}}$, satisfies $(nW\dot{m}^{\prime}_{\text{single}})_{\text{initial}}=(nW\dot{m}^{\prime}_{\text{single}})_{\text{optimal}}.$ This approach ensures that the dimensional designs deliver the same thermal power as the baseline configuration. The corresponding dimensional properties, including $L$, $D$, $t$, $W$, $n$, $\dot{m}^{\prime}_{\text{single}}$, and $\mathcal{D}$, are provided in the supplementary material (Table S2).
@@ 4 Design optimization considering additive manufacturing and fouling constraints
In cHEX design, thinner plates are generally preferred because they reduce axial heat conduction and minimize lateral conduction resistance, enhancing thermal performance. However, the minimum achievable plate thickness is limited by the capabilities of the selected manufacturing process and material. The goal of this section is to account for constraints related to manufacturing, fouling, and structural limitations when designing a cHEX for maximal compactness for a given pressure drop and effectiveness. To this end, we will introduce constraints on plate thickness and spacing into the optimal design problem. In the previous section, the nondimensional plate geometry was defined using a fixed ratio between plate thickness and spacing ($t^{\ast}/D^{\ast}$), which implicitly restricted the ability to impose a specific, manufacturable design plate thickness $t_{d}$. To explicitly introduce the minimal plate thickness, the original optimization problem in Eq. 15 is reformulated as follows:
$\max_{D^{\ast}>0,\ L^{\ast}>0,\ t^{\ast}>0}$ $\qquad Q^{\diamond}\left(D^{\ast},L^{\ast},t^{\ast}\right),$ (16)
s.t. $\qquad\varepsilon\left(D^{\ast},L^{\ast},t^{\ast};k_{w}/k\right)=\varepsilon_{d},$
$\qquad t^{\ast}=t_{d}/t_{\text{ref}},$
$\qquad D^{\ast}\geq D_{\text{min}}/t_{\text{ref}}.$
In addition to the plate thickness, a restriction on the plate spacing is imposed to reduce the risk of fouling. Specifically, the plate spacing $D$ must remain greater than a prescribed minimum value $D_{\text{min}}$, based on practical operating considerations. In the nondimensional formulation, this is enforced through the constraint $D^{\ast}\geq D_{\min}/t_{\text{ref}}$.
@@
The optimizations in this section use the same reference values, fluid properties, pressure drop, and effectiveness as the baseline heat exchanger, outlined in Section 3 and detailed in Table 1. The results of these optimizations are illustrated in Figure 2, which compares three distinct optimization strategies aimed at maximizing power density.
@@
The first strategy, labeled “Uniform Thickness Constraint”, applies a uniform $t_{d}$=0.5 mm across all materials, with no constraint on plate spacing. The second strategy, labeled “Uniform Thickness and Fouling Constraints”, introduces a fouling constraint that imposes $D_{min}$=0.8 mm, while still applying the same uniform plate thickness of 0.5 mm across all materials. The third strategy, labeled “Material-Specific Thickness and Fouling Constraints”, also incorporates both plate thickness and fouling constraints simultaneously. However, this approach applies a material-specific thickness constraint to account for variations in manufacturing resolution in AM between different materials and processes.
@@
To address AM resolutions for ceramic and metal heat exchangers, the plate thickness is determined based on the melt pool size, a characteristic of the manufacturing process. It is assumed that a sealed plate requires a thickness of twice the melt pool size. Consequently, the following plate thickness values are used: 0.5 mm for copper Jadhav2021 , 0.3 mm for aluminum Ghasemi2021 , and 0.25 mm for aluminum nitride (AlN), aluminum oxide (Al2O3Yao2023 , and austenitic steel Yang2021 . These values are characteristic of the Laser Powder Bed Fusion (LPBF) process, which is the most common technique for AM of metals and ceramics. For plastic heat exchangers, the plate thickness is not derived from melt pool size, as LPBF is not typically used for plastic printing. Instead, Material Extrusion (MEX), Vat Photopolymerization (VPP), and Sheet Lamination (SHL) processes are widely used Bhatia2023 , Ligon2017 . While these techniques can achieve thin walls with resolutions below 0.1 mm, a plate thickness of 0.1 mm is assumed in this study. This value was selected to balance manufacturing feasibility with structural integrity, as heat exchangers with wall thicknesses below 0.1 mm have been shown to be prone to plastic deformation Rua2015 .
In addition to the three optimization strategies, Figure 2 also presents nondimensional power density values for reference heat exchanger designs that are additive manufacturable. These are labeled as “Additive Manufacturable Reference” in the figure. In these designs, the plate spacing of the baseline heat exchanger is preserved, while the plate thickness is adjusted to reflect material-specific manufacturability constraints. To ensure comparability across materials, the channel length is modified as necessary so that each design achieves the same pressure drop of 170 Pa and the target effectiveness of 0.79. This definition aligns with the classical design methodology, where performance targets are met by varying channel length while maintaining the plate spacing Kays1984 . However, for copper, even with length adjustments, the design cannot attain the target effectiveness. The reason is that its asymptotic high-NTU effectiveness is limited to $\varepsilon_{\mathrm{NTU}\rightarrow\infty}=\frac{\textup{M}+1}{2\textup{M}+1}\approx 0.74,$ with $\textup{M}\approx 0.53$ obtained from Eq. 12, which is below the desired effectiveness of 0.79. Therefore, no Additive Manufacturable Reference design in copper is included in the figure. Consequently, the Additive Manufacturable Reference designs provide a consistent baseline for evaluating the effect of material properties on power density, while highlighting limitations in certain cases.
@@
Figure 2: Maximum nondimensional power densities obtained from optimizations with $\varepsilon_{d}~=~0.79$
@@
To provide additional context for the designs presented in Figure 2, Table 3 lists the nondimensional geometrical parameters and corresponding nondimensional power densities for each optimization strategy and the additive manufacturable reference designs. The geometrical values include nondimensional plate thickness $t^{\ast}$, plate spacing $D^{\ast}$, and heat exchanger length $L^{\ast}$, defined relative to the reference wall thickness $t_{\text{ref}}=$0.16\text{\,}\mathrm{mm}$\$. The dimensional equivalents are provided in Table S3 in the supplementary material.
Table 3: Nondimensional geometrical parameters and power densities for different optimization strategies and reference cases with $\varepsilon_{d}=0.79$. All geometrical values are scaled by the reference plate thickness $t_{\text{ref}}=$0.16\text{\,}\mathrm{mm}$\$.
Case Material $t^{\ast}$ $D^{\ast}$ $L^{\ast}$ $Q^{\diamond}\times 10^{6}$
Uniform Thickness Constraint Plastic 3.125 0.946 36.10 5.254
Austenitic steel 3.125 3.725 442.42 1.271
Ceramic (Al2O3) 3.125 4.278 565.46 1.090
Ceramic (AlN) 3.125 7.670 1890.17 0.386
Aluminum 3.125 8.413 2272.68 0.329
Copper 3.125 9.961 3197.05 0.244
Uniform Thickness and Fouling Constraints Plastic 3.125 5.00* 677.65 1.104
Austenitic steel 3.125 5.00* 684.29 1.083
Ceramic (Al2O3) 3.125 5.00* 708.41 1.010
Ceramic (AlN) 3.125 7.670 1890.17 0.386
Aluminum 3.125 8.413 2272.68 0.329
Copper 3.125 9.961 3197.05 0.244
Material-Specific Thickness and Fouling Constraints Plastic 0.625 5.00* 633.37 1.826
Austenitic steel 1.563 5.00* 652.01 1.477
Ceramic (Al2O3) 1.563 5.00* 662.81 1.429
Ceramic (AlN) 1.563 6.024 1177.34 0.685
Aluminum 1.875 7.175 1636.86 0.502
Copper 3.125 9.961 3197.05 0.244
Additive Manufacturable Reference Plastic 0.625 6.250 986.01 1.204
Austenitic steel 1.563 6.250 995.67 1.039
Ceramic (Al2O3) 1.563 6.250 1003.94 1.022
Ceramic (AlN) 1.563 6.250 1227.10 0.684
Aluminum 1.875 6.250 1469.77 0.458
@@
As shown in Figure 2, under the “Uniform Thickness Constraint” strategy, where the plate thickness is uniformly set to 0.5 mm ($t^{\ast}=3.125$) and the fouling constraint is not considered, plastic, the material with the lowest thermal conductivity, outperforms all other designs. The geometrical dimensions of this design are significantly smaller than those of the others (see Table 3). When the fouling constraint is introduced while still keeping the plate thickness uniform at 0.5 mm for all materials, the performance of plastic remains slightly better than that of austenitic steel and Al2O3. However, in the “Material-Specific Thickness and Fouling Constraints” strategy, where material-specific resolutions are explicitly accounted for, plastic again demonstrates a clearer performance advantage in power density due to its ability to utilize thinner plates.
In contrast, despite having the highest thermal conductivity among the materials considered, copper exhibits the lowest power density across all optimization strategies in Figure 2. This outcome highlights a broader trend observed in the results: heat exchangers made from high-conductivity materials such as copper, aluminum, and AlN consistently require increased lengths and plate spacings to achieve the target effectiveness of 0.79. As shown in Table 3, the fouling constraint remains inactive for these materials in all strategies, since their optimal spacings are naturally above the minimum threshold. However, their plate spacings are slightly smaller when material-specific thicknesses are used, compared to the uniform-thickness cases. In general, designs with thicker plates and high-conductivity materials result in larger heat exchangers, as longer channels and wider gaps are required to offset axial wall conduction and maintain the target effectiveness.